• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 07:55:39 on November 8,2009

    Entry number 298298

    keyword=Owl shift summary

    12:00 Shift started.
    1:30 Magnets aredown.
    2:00 Short controlled access, Rupesh went to hall and recycle the magnets.
    2:20 Jack is here, he went to hall to recycle the magnets.
    3:20 Left arm dipole is down for night. Right arm dipoleis ramping down, lets see if we can ramp it up tonight.
    3:45 Rupesh is working on dithering.
    4:38 Q3 is down in right arm. Target is Hydrogen for higher rates of oversampling.
    4:43 Q1 in right arm is down.
    7:47 Rupesh finished the oversampling studies.